The utility function for a consumer utility is U=30Q11/2Q21/2 . If the price per unit of Q1 is Kshs 10 and Kshs 5per unit of Q2, determine quantities Q1 and Q2 that the consumer should have to maximize utility if the consumer budgeted Kshs 350


1
Expert's answer
2022-02-09T10:54:29-0500

determine quantities Q1 and Q2 that the consumer should have to maximize utility if the consumer budgeted Kshs 350


We will determine the marginal utility functions for Q1 and Q2 as:

MU(Q1) = "\dfrac{∂U}{∂Q1}" = "\dfrac{15Q2^{\smash{0.5}}}{Q1^{\smash{0.5}}}"


MU(Q2) = "\dfrac{∂U}{∂Q2}" = "\dfrac{15Q1^{\smash{0.5}}}{Q2^{\smash{0.5}}}"

The condition for optimal consumer choice:

"\dfrac{MU(Q1)}{MU(Q2)}" = "\dfrac{P1}{P2}"

Therefore, "\dfrac{15Q2^{\smash{0.5}}}{Q1^{\smash{0.5}}}" ÷ "\dfrac{15Q1^{\smash{0.5}}}{Q2^{\smash{0.5}}}" = "\dfrac{10}{2}" "\dfrac{Q1}{Q2} = 2 ⇒ Q2 = 2Q1"


From the budget constraint:

10Q1+5Q2=350

But Q2=2Q1

Therefore 10Q1+5(2Q1) =350

⇒20Q1=350 

Therefore; Q1=17.5 and Q2=35 are quantities Q1 and Q2 that the consumer would have maximized utility.